Given the following information, calculate the straight-line depreciation of the equipment. Equipment had a cash purchase price of $50,000; sales tax of $3,000; and installation costs of $1,000. Salvage value is expected to be $2,000. The useful life is expected to be 20 years.
Answer #1

Total cost of Equipment = 50,000+3,000+1,000 = $54,000

Salvage Value = $2,000

Straight line depreciation (Cost - Salvage Value)/Life

= (54,000-2,000)/20

= $2,600